a 4 poles generator has a lap wound armature with 50 slots and with 16 conductors per slot. The useful flux per pole is 30 milli Weber.Determine the speed assuming the machine must be driven to generate an e.m.f of 240 volts.


1
Expert's answer
2021-02-17T07:04:47-0500

"E = \frac{\phi ZN \times P}{60\times A}"

"E = \frac{240\times 60\times 4}{3\times 10^-3\times 800\times 4}"

=600 rpm